OpenStudy (anonymous):

Find the x- and y-intercepts of y= -3x-9

OpenStudy (anonymous):

kgw22, you need to learn how to do these, not just get answers. set x=0 and solve to find your y-intercept. set y=0 and solve to find your x-intercept. it's that easy, try it!

OpenStudy (anonymous):

idk how to do it

OpenStudy (anonymous):

to find the intercepts replace one of the variables with 0 and solve for the other variable. it will be the intercept i.e. y=-3(0)-9 -> y=9 is the y intercept 0= -3x-9 -> x=-3 is the x intercept

OpenStudy (anonymous):

I think you are doing your home work this is not a place to only get answers more than understanding

OpenStudy (anonymous):

i just told you how. so lets plug in the values. lets start with the y-intercept. so the line will intersect the y-axis when x=0 because that's where the y-axis is located.

OpenStudy (anonymous):

This guy is either mentally disabled or trolling. Don't waste your time.

OpenStudy (anonymous):

so we plug in 0 for x and see what we get: y=-3(0)-9 y=0-9 y=-9 so our point is (x,y) and since we chose x to be 0 the point is (0,-9)

OpenStudy (anonymous):

now try the next one